OpenStudy (anonymous):

x−6x^(1/3) domain pls

OpenStudy (anonymous):

R

OpenStudy (anonymous):

All real number R

OpenStudy (anonymous):

oh wow. what about the x and y-int

OpenStudy (anonymous):

I dislike this kind of question, domain is part of function specification.

OpenStudy (anonymous):

yeah me too.

OpenStudy (anonymous):

can you help me with the x and y-int of this pls

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

when you set x = 0 , you will get y = 0 as well when you set y 0, you will get this equation which is x−6x^(1/3) = 0 To solve it: x = 6 x^(1/3) cube both sides x^3 = 216 x x^3 - 216x = 0 x (x^2 - 216) = 0 x = 0 OR x^2 -216=0

OpenStudy (anonymous):

my software didnt accept the ans

OpenStudy (anonymous):

hmmmmmmmm

OpenStudy (anonymous):

Is this an equation or what?

OpenStudy (anonymous):

@Jinnie

OpenStudy (anonymous):

its a function f(x)= x−6x^(1/3)

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

Can you try (0,0) and (14.69 , 0)

OpenStudy (anonymous):

neither worked

OpenStudy (anonymous):

So I don't know because it's complicated equation Sorry to say that

OpenStudy (anonymous):

ok what about this

OpenStudy (anonymous):

?

OpenStudy (anonymous):

(x)=3cosx−cos^(3)x on the interval 0<x<2π

OpenStudy (anonymous):

whats the x-int for that

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

Is it f(x) in the begining or just (x)

OpenStudy (anonymous):

f(x)=

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

ok np

OpenStudy (anonymous):

it's gonna be pi/2

OpenStudy (anonymous):

(pi/2 , 0)

OpenStudy (anonymous):

right? @Jinnie

OpenStudy (anonymous):

program says no

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

oh sorry it's (pi/2 , 0) OR (3pi/2 , 0)

OpenStudy (anonymous):

check these please @Jinnie

OpenStudy (anonymous):

look at the doc attached

OpenStudy (anonymous):

you wrote it in wrong order and what is the union sign between them?

OpenStudy (anonymous):

Don't make union sigh between them it's just a points

OpenStudy (anonymous):

i tried (pi/2,0) and (3pi/2,0) neither worked

OpenStudy (anonymous):

I don't really know :( I'm sorry about that

OpenStudy (anonymous):

ok might you know the local maximum?

OpenStudy (anonymous):

Actually I can't remember it I'm sorry again

OpenStudy (anonymous):

ok what about this -9sin^2(x) cos(x) can you tell which interval this is increasing and decreasing on

OpenStudy (anonymous):

not increasing and decreasing but concave up and down

OpenStudy (anonymous):

I think you have to take the derivative and then make it equal to 0 and find the values for x

OpenStudy (anonymous):

Can you do that?

OpenStudy (anonymous):

ive tried doing that but im not getting the right answer

OpenStudy (anonymous):

-9sin^2(x) cos(x)<<<< this is already in derivative form so -9sin^2(x) cos(x)= 0 and i get nonsense

OpenStudy (anonymous):

I don't really remember this stuff I'm sorry @Jinnie

OpenStudy (anonymous):

ok np

OpenStudy (anonymous):

wow this is a long thread. what is the actual question? domain is all real numbers. after than i am lost as to what the question is

OpenStudy (anonymous):

there are a couple but can we start with this pls -9sin^2(x) cos(x)....where is it concave up and down....its already in second derivative form

OpenStudy (anonymous):

\[f(x)=(x)=3\cos(x)−\cos^3(x)\]

OpenStudy (anonymous):

what is the original function?

OpenStudy (anonymous):

3cosx−cos^(3)x for 0<x<2π

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

so first derivative is \[-3\sin(x)+3\cos^2(x)\sin(x)\] right?

OpenStudy (anonymous):

-3sin^3 x is the first derivative

OpenStudy (anonymous):

so it is

OpenStudy (anonymous):

and therefore your function is increasing on \((\pi, 2\pi)\) and decreasing on \((0,\pi)\)

OpenStudy (anonymous):

as the cube doesn't change the sign, but the -3 out front does. so far so good?

OpenStudy (anonymous):

so far so good

OpenStudy (anonymous):

second derivative \[f''(x)=-9\sin^2(x)\cos(x)\]

OpenStudy (anonymous):

You are so intelligent @satellite73

OpenStudy (anonymous):

now we only need to worry about \[\cos(x)\] because \(\sin^2(x)\geq 0\) on the interval \((\frac{\pi}{2},\frac{3\pi}{2})\) we have \(\cos(x)<0\)and so \(-\cos(x)>0\) and that is where your original funcion is concave up

OpenStudy (anonymous):

it will be concave down on \((0,\frac{\pi}{2})\) and also on \((\frac{3\pi}{2},2\pi)\)

OpenStudy (anonymous):

yea that works thanks

OpenStudy (anonymous):

that about takes care of it. not sure what "works" means, but i guess it is on line homework

OpenStudy (anonymous):

its an online hw yea...do you know the point of inflection